# Neural Beam Field for Spatial Beam RSRP Prediction

### v1

@@ Abstract
Accurately predicting beam-level reference signal received power (RSRP) is essential for beam management in dense multi-user wireless networks, yet challenging due to high measurement overhead and fast channel variations. This paper proposes Neural Beam Field (NBF), a hybrid neural-physical framework for efficient and interpretable spatial beam RSRP prediction. Central to our approach is the introduction of the Multi-path Conditional Power Profile (MCPP), which bridges site-specific multipath propagation with antenna/beam configurations via closed-form analytical modeling. We adopt a decoupled “blackbox-whitebox” design: a Transformer-based deep neural network (DNN) learns the MCPP from sparse user measurements and positions, while a physics-inspired module analytically infers beam RSRP statistics. To improve convergence and adaptivity, we further introduce a Pretrain-and-Calibrate (PaC) strategy that leverages ray-tracing priors and on-site calibration using RSRP data. Extensive simulations results demonstrate that NBF significantly outperforms conventional table-based channel knowledge maps (CKMs) and pure blackbox DNNs in prediction accuracy, training efficiency, and generalization, while maintaining a compact model size. The proposed framework offers a scalable and physically grounded solution for intelligent beam management in next-generation dense wireless networks.
Index Terms:
Neural Beam Field, Multipath Conditional Power Profile, Channel Knowledge Map, Beam‐level RSRP, Transformer
@@ I Introduction
With the continuous densification of mobile communication networks, network operators face unprecedented challenges in maintaining scalable, highly spectral-efficient and reliable service coverage. Although massive Multiple-Input Multiple-Output (MIMO) and millimeter-wave technologies promise significant improvements in capacity and user throughput, the practical deployment of dense wireless networks is significantly constrained by the difficulty of timely acquiring Channel State Information (CSI) under realistic multi-user scenarios[1]. Traditionally, accurate CSI acquisition involves comprehensive measurements at high spatial-temporal-spectral resolutions, which inevitably leads to prohibitive overheads and excessive latency, severely limiting network scalability and real-time applicability [2, 3].
Traditional channel estimation for MIMO beamforming mainly relies on two methods, i.e., pilot-based CSI estimation and beam scanning. The former method provides full CSI but incurs high overhead and latency, with the pilots given in the fifth-generation (5G) standards in the form of downlink CSI-RS or uplink SRS signals[2, 3]. In comparison, the beam scanning method selects beam pairs from a codebook to estimate Reference Signal Received Power (RSRP) or Signal-to-Interference-plus-Noise Ratio (SINR) for user-specific beam pair links (BPLs). To alleviate the burden of CSI acquisition, reduced-dimensional channel indicators such as RSRP or SINR have been widely adopted in practical network deployments [4, 5]. Nevertheless, even these simplified RSRP-based measurements can hardly satisfy real-time requirements due to varying propagation conditions and user mobility. Moreover, dense network conditions exaggerate the complexity, as complete measurement of RSRP for all user-beam combinations often exceeds channel coherence time, rendering CSI outdated before it can be utilized for resource allocation or beam scheduling [6]. Moreover, the beam-specific RSRP is highly directional and thus relies on the site-specific blockages/shadowing/multi-path conditions. These critical issues demand innovative approaches capable of effectively predicting spatial RSRP fields under diverse beamforming configurations and environmental conditions.
@@
  • We introduce the new notion of Multi-path Conditional Power Profile (MCPP), which serves as an essential intermediary between physical wave propagation in a site-specific environment and the add-on antenna panel/beam configurations. Based on MCPP, we further derive analytical formulas for beam RSRP statistics, which facilitate easier learning and generalization under diverse settings.
  • We propose a decoupled “blackbox-whitebox” approach to construct NBF, by 1) learning the site-specific/irregular MCPPs using a tailored design of Transformer-based DNN, and 2) inferring the beam RSRP statistics based on analytical formulas. This “whitebox” analytical modeling combined with “blackbox” neural representations results in an interpretable and accurate predictive framework.
  • To achieve better convergence and optimality for the highly nonlinear mapping task in NBF, we further propose a Pretrain-and-Calibrate (PaC) strategy in cases when prior (but not necessarily exact) information about MCPPs are available. Leveraging on tailored MCPP loss functions, the PaC strategy helps tame the NBF into a desired locally optimal point, which could then adapt to unknown environmental factors by on-site RSRP measurements.
Numerical results validate the analytical formulas through extensive Monte Carlo (MC) simulations. Moreover, it is shown that the proposed approach significantly outperforms conventional table-based CKMs and pure black-box neural models in terms of prediction accuracy, storage efficiency, and generalization capability. Our proposed NBF method for spatial beam RSRP prediction holds potential in intelligent beam management and user scheduling in dense wireless networks.
@@ II-A Antenna Panel Rotation and Array Response Vector
Denote $\boldsymbol{u}_{\text{tx}}\triangleq[\cos\varphi_{\text{tx}}\sin\theta_{\text{tx}},\sin\varphi_{\text{tx}}\sin\theta_{\text{tx}},\cos\theta_{\text{tx}}]^{T}$ as the 3-dimensional (D) unit direction vector of a given wave propagation path w.r.t. the origin, with $\Pi_{\text{tx}}\triangleq(\varphi_{\text{tx}},\theta_{\text{tx}})$ denoting the corresponding (horizontal, vertical) direction of departure (DOD) angle in the GCS. Consider one typical antenna panel which locates in the Y-O-Z plane in the GCS. The components of its array response vector (ARV) along the $y\text{-}$ and $z\text{-}$axis are given by
@@ II-B Channel Model
where $(\boldsymbol{u}_{\text{tx},l},\,\boldsymbol{u}_{\text{rx},l},\,\tau_{l}\bigr{)}$ collectively represents the seven-dimensional (7-D) path condition. Note that we choose the unit direction vectors $\boldsymbol{u}_{\text{tx},l}$ and $\boldsymbol{u}_{\text{rx},l}$ instead of the angles $\Pi_{\text{tx},l}$ and $\Pi_{\text{rx},l}$ for better representation and easier learning of periodic angle features. Furthermore, we introduce the new notion of Multi-path Conditional Power Profile (MCPP) for all paths in $\mathcal{L}$, denoted as
@@ II-D Practical Collection/Synthetic Generation of RSRP Data
On the other hand, as a preliminary study, we also provide ways for synthetic generation of RSRP data. In compliance with the 3GPP TR38.901 hybrid channel model[20], we employ SionnaRT[21] to generate site-specific MCPPs for a given city region based on ray tracing, while utilizing QuaDRiGa[22] to generate spatially consistent random perturbations for the MCPPs to simulate unknown environmental factors. The final MCPP for RSRP generation at each location is then obtained as the weighted sum of the above deterministic and random MCPP components.
@@ III-A Neural Beam Field Design
The overall NBF design is illustrated in Fig. 2, which consists of 1) a Transformer-based blackbox model $f_{\vartheta}(\mathbf{x})$ that predicts the MCPP at a given UE location $\mathbf{x}$, and 2) a whitebox model that maps to beam RSRP statistics based on Proposition 1. Combined together, the NBF function is given by
$\mathrm{NBF}(\mathbf{x},\mathbf{w}_{\text{tx}})\triangleq\big{\{}\mu_{\mathrm{RSRP}}\big{(}f_{\vartheta}(\mathbf{x}),\mathbf{w}_{\text{tx}}\big{)},\,\sigma_{\mathrm{RSRP}}^{2}\big{(}f_{\vartheta}(\mathbf{x}),\mathbf{w}_{\text{tx}}\big{)}\big{\}}.$ (15)
In the following, we focus on the DNN design.
Firstly, for the input of UE location $\mathbf{x}$, we adopt the random Fourier position embedding[24] to obtain a 256-D vector $\mathbf{X}_{\text{UE}}\in\mathbb{R}^{1\times 256}$, which can then be treated as a token for subsequent Transformer blocks. In addition, in order to output the MCPP which consists of $L$ SCPs, we also place $L$ learnable target tokens $\mathbf{X}_{\text{target}}\in\mathbb{R}^{L\times 256}$ at the input. Secondly, the concatenated input tokens $\mathbf{X}\triangleq[\mathbf{X}_{\text{UE}};\mathbf{X}_{\text{target}}]$ are processed by the Transformer encoder, comprising $N$ blocks of multi-head self-attention (MSA), MLP, layer normalization (LN), and residual connections[23]. Notably, the learnable target tokens attend to the UE location token along the blocks to extract key information related to location-specific MCPPs. Finally, each of the target tokens processed by Transformer is fed to two MLPs, i.e., 1) a two-layer regression head that produces 8 features corresponding to one SCP in (6), thereby composing the predicted MCPP $\hat{\mathbf{Y}}=f_{\vartheta}(\mathbf{x})$; and 2) a one-layer classification head with output $a_{l}\in\{0,1\}$ that tells the existence of path $l$ (in cases with fewer than $L$ paths).
@@ III-B End-to-End Learning Based on RSRP Measurements
Based on the above NBF design, we then train the DNN such that the end-to-end (blackbox-whitebox) prediction of beam RSRP statistics matches as close as possible to the corresponding labels from measurement data. Regression loss functions such as the smooth L1 loss or mean absolute error (MAE) loss could be employed.
@@ III-C MCPP-Prior Aided Pretraining and On-Site Calibration
In the pretraining phase, MCPP labels in the form of (7) are available, denoted as $\bar{\mathbf{Y}}$. We aim to minimize the difference between the predicted MCPP $\hat{\mathbf{Y}}$ and ground-truth MCPP $\bar{\mathbf{Y}}$. However, two issues exist, i.e., 1) some labels $\bar{\mathbf{Y}}$ might have paths fewer or more than $L$, and 2) the permutation-invariant path indices require per-path matching between $\hat{\mathbf{Y}}$ and $\bar{\mathbf{Y}}$. Regarding the first issue, note that $L$ could be chosen as the typical (maximum) number of significant paths under given environment type and carrier frequency. In some outlier locations with more than $L$ paths, clustering methods such as weighted K-means could be employed to merge them into $L$ significant paths. On the other hand, for locations with fewer than $L$ paths, we utilize the path existence prediction $a_{l}$ along with binary cross-entropy (BCE) loss to learn the situation. Regarding the second issue, we employ the Hungarian algorithm for one-to-one path matching. The above processing thus allows us to capture the difference between $\hat{\mathbf{Y}}$ and $\bar{\mathbf{Y}}$, using regression loss functions such as MAE. The overall loss function is then taken as the weighted sum of the MCPP regression loss and path existence BCE loss, whereby the details are omitted for brevity.
Finally, in the calibration phase, the on-site RSRP measurements could be utilized to fine-tune the NBF based on the process in Section III-B. In summary, the PaC strategy helps tame the NBF into a desired locally optimal point, which could then adapt quickly to other unknown environmental factors implicitly manifested in the on-site RSRP measurements.
@@ IV Numerical Results
This section presents numerical results that validate the analytical formulas and evaluate the effectiveness of NBF. First, we verify the analytical formulas from Proposition 1 through extensive MC simulations. Second, we assess the performance of NBF by comparing it with two benchmark schemes, including a pure blackbox model based on MLP, and a table-based CKM with fixed interpolation rules using IDW. The MLP baseline performs end-to-end regression of user performance, using normalized user coordinates and sin-cos encoded DFT beam spatial frequencies as inputs. Its model size is slightly larger than that of the NBF to ensure a fair comparison of model architectures. The IDW baseline interpolates the MCPP at a target location from nearby samples and plugs the estimate into the white-box analytical expression to obtain user performance. The following parameters are used unless otherwise specified: $h_{\text{BS}}=20$ m, $h_{\text{UE}}=1.5$ m, $L=10$, $f_{c}=3.5$ GHz, antenna panel dimensions $N_{\mathrm{tx,y}}=8$ and $N_{\mathrm{tx,z}}=4$, a square area with side length 256 m, a panel down-tilt angle of $15^{\circ}$, peak antenna element gain of 8 dBi, and half-wavelength inter-element spacing.

### v2

@@ Abstract
Accurately predicting beam-level reference signal received power (RSRP) is essential for beam management in dense multi-user wireless networks, yet challenging due to high measurement overhead and fast channel variations. This paper proposes Neural Beam Field (NBF), a hybrid neural-physical framework for efficient and interpretable spatial beam RSRP prediction. Central to our approach is the introduction of the Multi-path Conditional Power Profile (MCPP), a learnable physical intermediary representing the site-specific propagation environment. This approach decouples the environment from specific antenna/beam configurations, which helps the model learn site-specific multipath features and enhances its generalization capability. We adopt a decoupled “blackbox-whitebox” design: a Transformer-based deep neural network (DNN) learns the MCPP from sparse user measurements and positions, while a physics-inspired module analytically infers beam RSRP statistics. To improve convergence and adaptivity, we further introduce a Pretrain-and-Calibrate (PaC) strategy that leverages ray-tracing priors for physics-grounded pretraining and then RSRP data for on-site calibration. Extensive simulation results demonstrate that NBF significantly outperforms conventional table-based channel knowledge maps (CKMs) and pure blackbox DNNs in prediction accuracy, training efficiency, and generalization, while maintaining a compact model size. The proposed framework offers a scalable and physically grounded solution for intelligent beam management in next-generation dense wireless networks.
@@ I Introduction
With the continuous densification of mobile communication networks, network operators face unprecedented challenges in maintaining scalable, highly spectral-efficient and reliable service coverage. Although massive Multiple-Input Multiple-Output (MIMO) and millimeter-wave technologies promise significant improvements in capacity and user throughput, the practical deployment of dense wireless networks is significantly constrained by the difficulty of timely acquiring Channel State Information (CSI) under realistic multi-user scenarios[1]. Traditionally, accurate CSI acquisition involves comprehensive measurements at high spatial-temporal-spectral resolutions, which inevitably leads to prohibitive overheads and excessive latency, severely limiting network scalability and real-time applicability [2].
Traditional channel estimation for MIMO beamforming mainly relies on two methods, i.e., pilot-based CSI estimation and beam scanning. The former method provides full CSI but incurs high overhead and latency, with the pilots given in the fifth-generation (5G) standards in the form of downlink CSI-RS or uplink SRS signals[2]. In comparison, the beam scanning method selects beam pairs from a codebook to estimate Reference Signal Received Power (RSRP) or Signal-to-Interference-plus-Noise Ratio (SINR) for user-specific beam pair links (BPLs). To alleviate the burden of CSI acquisition, reduced-dimensional channel indicators such as RSRP or SINR have been widely adopted in practical network deployments [3]. Nevertheless, even these simplified RSRP-based measurements can hardly satisfy real-time requirements due to varying propagation conditions and user mobility. Moreover, dense network conditions exaggerate the complexity, as complete measurement of RSRP for all user-beam combinations often exceeds channel coherence time, rendering CSI outdated before it can be utilized for resource allocation or beam scheduling [4]. Moreover, the beam-specific RSRP is highly directional and thus relies on the site-specific blockages/shadowing/multi-path conditions. These critical issues demand innovative approaches capable of predicting spatial RSRP fields under diverse beamforming configurations and environmental conditions.
@@
  • First, we introduce the new notion of Multi-path Conditional Power Profile (MCPP), which serves as an essential intermediary between physical wave propagation in a site-specific environment and the add-on antenna panel/beam configurations. Based on MCPP, we further derive analytical formulas for beam RSRP statistics, which facilitate easier learning and generalization under diverse settings. Therefore, the incorporation of beam RSRP and MCPP-based physical grounding provides one key innovation compared with existing neural field methods.
  • Second, we propose a decoupled “blackbox-whitebox” approach to construct NBF, by 1) learning the site-specific MCPPs using a tailored design of Transformer-based DNN, and 2) inferring the beam RSRP statistics based on analytical formulas. This “whitebox” analytical modeling combined with “blackbox” neural representations results in an interpretable and accurate predictive framework.
  • Finally, we further propose a Pretrain-and-Calibrate (PaC) strategy in cases when prior information about MCPPs are available. Leveraging on 1) tailored MCPP loss functions that could handle un-ordered/varying number of paths and 2) regularized calibration loss that preserves the model’s physical groundings and enhances finetuning stability, the PaC strategy helps tame the NBF into a robust locally optimal point that could then adapt to unknown/random factors by on-site RSRP measurements.
Numerical results validate the analytical formulas through extensive Monte Carlo (MC) simulations. Moreover, it is shown that the proposed approach significantly outperforms conventional table-based CKMs and pure blackbox neural models in terms of prediction accuracy, storage efficiency, and generalization capability. Our proposed NBF method for spatial beam RSRP prediction holds potential in intelligent beam management and user scheduling in dense wireless networks.
@@ II-A Antenna Panel Rotation and Array Response Vector
Denote $\boldsymbol{u}_{\text{tx}}\triangleq[\cos\varphi_{\text{tx}}\sin\theta_{\text{tx}},\sin\varphi_{\text{tx}}\sin\theta_{\text{tx}},\cos\theta_{\text{tx}}]^{T}$ as the 3-dimensional (D) unit vector of a given wave propagation path w.r.t. the origin, with $\Pi_{\text{tx}}\triangleq(\varphi_{\text{tx}},\theta_{\text{tx}})$ denoting the corresponding (horizontal, vertical) direction of departure (DOD) angle in the GCS. Consider one typical antenna panel which locates in the Y-O-Z plane in the GCS. The components of its array response vector (ARV) along the $y\text{-}$ and $z\text{-}$axis are given by
@@ II-B Channel Model
where $(\boldsymbol{u}_{\text{tx},l},\,\boldsymbol{u}_{\text{rx},l},\,\tau_{l}\bigr)$ collectively represents the 7-D path condition. Note that we choose the unit direction vectors $\boldsymbol{u}_{\text{tx},l}$ and $\boldsymbol{u}_{\text{rx},l}$ instead of the angles $\Pi_{\text{tx},l}$ and $\Pi_{\text{rx},l}$ for better representation and easier learning of periodic angle features. Furthermore, we introduce the new notion of Multi-path Conditional Power Profile (MCPP) for all paths in $\mathcal{L}$, denoted as
@@ II-D Practical Collection/Synthetic Generation of RSRP Data
On the other hand, as a preliminary study, we also provide ways for synthetic generation of RSRP data. In compliance with the 3GPP TR38.901 hybrid channel model[18], we employ SionnaRT[19] to generate site-specific MCPPs for a given city region based on ray tracing, while utilizing QuaDRiGa[20] to generate spatially consistent random perturbations for the MCPPs to simulate unknown/random environmental factors. The final MCPP at each location is then obtained as the weighted sum of the above deterministic and random MCPP components, with normalized weight $0\leq\beta<1$ for the latter.
@@ III-A Neural Beam Field Design
The overall NBF design is illustrated in Fig. 2, which consists of 1) a Transformer-based blackbox model $f_{\vartheta}(\mathbf{x})$ that predicts the MCPP at a given UE location $\mathbf{x}$, and 2) a whitebox model that maps to beam RSRP statistics based on Proposition 1. For the purpose of exposition, we consider the mean RSRP as the output while higher order statistics could be treated in similar manners. Therefore, the NBF function is given by
$\mathrm{NBF}(\mathbf{x},\mathbf{w}_{\text{tx}})\triangleq\mu_{\mathrm{RSRP}}\big(f_{\vartheta}(\mathbf{x}),\mathbf{w}_{\text{tx}}\big),$ (15)
with UE location $\mathbf{x}$ and beam configuration $\mathbf{w}_{\text{tx}}$ as the input. In the following, we focus on the DNN design.
Firstly, for the input of UE location $\mathbf{x}=(x,y)$, we adopt the random Fourier position embedding[22] to obtain a $d_{\text{model}}$-D feature vector $\mathbf{X}_{\text{UE}}\in\mathbb{R}^{1\times d_{\text{model}}}$, which can then be treated as a token for subsequent Transformer blocks. In addition, in order to output the MCPP which consists of $L$ SCPs, we also place $L$ learnable target tokens $\mathbf{X}_{\text{target}}\in\mathbb{R}^{L\times d_{\text{model}}}$ at the input. Secondly, the concatenated input tokens $\mathbf{X}\triangleq[\mathbf{X}_{\text{UE}};\mathbf{X}_{\text{target}}]$ are processed by the Transformer encoder, comprising $N$ blocks of multi-head self-attention (MSA), multi-layer perceptron (MLP), layer normalization (LN), and residual connections[21]. Notably, the learnable target tokens attend to the UE location token along the blocks to extract key information related to location-specific MCPPs. Finally, each of the target tokens processed by Transformer is fed to two MLPs, i.e., 1) a two-layer regression head that produces 8 features corresponding to one SCP in (6), thereby composing the predicted MCPP $\hat{\mathbf{Y}}=f_{\vartheta}(\mathbf{x})$; and 2) a one-layer classification head with output $a_{l}\in\{0,1\}$ that tells the existence of path $l$ (in cases with fewer than $L$ paths).
@@ III-B End-to-End Learning Based on RSRP Measurements
Based on the above NBF design, we then train the DNN such that the end-to-end (blackbox-whitebox) prediction of beam-specific mean RSRP matches as close as possible to the corresponding labels from measurement data. Note that the whitebox module, which implements the analytical formulas from Proposition 1, contains no trainable parameters. It serves as a fixed, differentiable mapping function that connects the MCPP to the mean RSRP. This structure allows for end-to-end gradient-based optimization of the blackbox part. Finally, the Smooth-$L_{1}$ function is chosen as the regression loss $\mathcal{L}_{\text{RSRP}}$ because it is less sensitive to outliers than the $L_{2}$ loss. The gradients from this loss back-propagate through the whitebox module to update only the blackbox parameters $\vartheta$.
@@ III-C MCPP-Prior Aided Pretraining and On-Site Calibration
In the pretraining phase, MCPP labels in the form of (7) are available, denoted as $\bar{\mathbf{Y}}$. We aim to minimize the difference between the predicted MCPP $\hat{\mathbf{Y}}$ and ground-truth MCPP $\bar{\mathbf{Y}}$. However, two issues exist, i.e., 1) some labels $\bar{\mathbf{Y}}$ might have paths fewer or more than $L$, and 2) the permutation-invariant path indices require per-path matching between $\hat{\mathbf{Y}}$ and $\bar{\mathbf{Y}}$. Regarding the first issue, note that $L$ could be chosen as the typical (maximum) number of significant paths under given environment type and carrier frequency. In some outlier locations with more than $L$ paths, clustering methods such as weighted K-means could be employed to merge them into $L$ significant paths. On the other hand, for locations with fewer than $L$ paths, we utilize the path existence prediction $a_{l}$ along with binary cross-entropy (BCE) loss to learn the situation.
To handle un-ordered/varying number of paths, we propose a set prediction loss based on bipartite matching. For each training sample, we first find an optimal one-to-one assignment between the $L$ predicted paths and the ground-truth paths using the Hungarian algorithm. The assignment cost matrix considers both the regression difference of path parameters and the classification score for path existence. After matching, the total loss for the pre-training stage, $\mathcal{L}_{\text{pretrain}}$, is the weighted sum of a classification loss $\mathcal{L}_{\text{cls}}$ and a regression loss $\mathcal{L}_{\text{reg}}$, i.e.,
$\mathcal{L}_{\text{pretrain}}\triangleq\mathcal{L}_{\text{cls}}+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}},$ (16)
where $\mathcal{L}_{\text{cls}}$ adopts a binary cross-entropy (BCE) loss computed over all $L$ predicted path existence logits, and $\mathcal{L}_{\text{reg}}$ adopts the Smooth-$L_{1}$ loss computed only between the parameters of the matched pairs, with the weight $\lambda_{\text{reg}}$ to balance the two losses.
During the calibration phase, besides the loss $\mathcal{L}_{\text{RSRP}}$ for mean RSRP prediction, to enhance the stability of fine-tuning and mitigate potential catastrophic forgetting, we introduce a regularization loss $\lambda_{\text{feat}}$ with weight $\lambda_{\text{feat}}$ based on feature representation alignment, resulting in a joint loss function, i.e.,
$\mathcal{L}_{\text{finetune}}\triangleq\mathcal{L}_{\text{RSRP}}+\lambda_{\text{feat}}\mathcal{L}_{\text{feat}}.$ (17)
More specifically, the parameters $\vartheta$ of the pretrained network are loaded and frozen to serve as a reference model. The term $\mathcal{L}_{\text{feat}}$ then quantifies the mean squared error (MSE) between the intermediate feature representations produced by the current model and those produced by the reference model. This regularized calibration approach effectively transfers the inductive bias acquired during pretraining to the fine-tuning stage, thus preserving the model’s physical groundings while efficiently adapting to the unknown environmental factors.
III-D Computational Complexity Analysis
During forward inference, the NBF invokes a blackbox DNN with $L+1$ tokens passed through $N$ Transformer blocks to predict $L$ SCPs, which are then fed to the whitebox to map into mean RSRP. According to [23], the inference complexity is $O\big(N\cdot[(L+1)^{2}\cdot d_{\text{model}}+(L+1)\cdot d_{\text{model}}^{2}]\big)$, with $d_{\text{model}}$ being the number of dimensions in each token. In addition, the whitebox module processes each SCP independently and thus is of complexity $O(L)$. In comparison, the training stage involves the same complexity order as in the inference stage[24], with the additional cost for loss evaluation. In particular, the MCPP path matching adopts the classic Hungarian algorithm which is of complexity $O(L^{3})$. Finally, we adopt the Adam optimizer which is guaranteed to converge to at least one locally optimal point[25]. Also note that, whether training or inference, the computational complexity of NBF is independent of the geographical area size, which affects only the value range of UE locations and the number of available samples.
@@ IV Numerical Results
In this section, we first verify Proposition 1 through extensive MC simulations, and then demonstrate the effectiveness of NBF compared with representative benchmark schemes. As discussed in Section I, RSRP prediction based on spatially distributed RSRP measurements mainly includes classic spatial interpolation methods[7, 8, 9, 10, 11, 12, 13] and DNN-based methods[14, 15, 16], which have yet to explicitly incorporate beam RSRP. As a preliminary study, we consider inverse distance weighting (IDW) as a representative for the former and an MLP-based DNN for the latter, while leaving more exhaustive comparisons for future work. IDW performs spatial interpolation from nearby grids [^2] The neighbor range is empirically chosen as $3d$ for desired performance., which consists of two variants, i.e., 1) interpolating the mean RSRP labels directly, and 2) interpolating the MCPP first and then use the whitebox module for inference. The MLP baseline performs end-to-end regression of mean RSRP, using normalized user coordinates and sin-cos encoded DFT beam spatial frequencies as inputs. Its model size is slightly larger than that of the NBF to ensure a fair comparison. The following parameters are used: $h_{\text{BS}}=20$ m, $h_{\text{UE}}=1.5$ m, $L=10$, $f_{c}=3.5$ GHz, $N_{\mathrm{tx,y}}=8$, $N_{\mathrm{tx,z}}=4$, $\beta=0.5$, $\lambda_{\text{reg}}=5$, $\lambda_{\text{feat}}=0.1$, $d_{\text{model}}=256$, a square area with side length 256 m and $d=1$ m, a panel down-tilt angle of $15^{\circ}$, peak antenna element gain of 8 dBi, and half-wavelength inter-element spacing.